Show that the function f given by f(x)=|x| is continuous on R^n. [Hint: Consider |x-a|^2=(x-a).(x-a).]
1
Expert's answer
2020-11-26T10:40:09-0500

Let "\epsilon>0" be given. We want to find a "\delta=\delta(\epsilon) >0" such that for any "x_0 \in \mathbb{R}" if "|x-x_0|<\delta" then, "|f(x)-f(x_0)|<\epsilon"

Now,


"|f(x)-f(x_0)|=||x|-|x_0|| \leq|x-x_0|<\delta=\epsilon" if "\delta=\epsilon."


"\implies |f(x)-f(x_0)|<\epsilon"

Hence, "f(x)=|x|" is continuous on "\mathbb{R}"